OpenStudy (anonymous):

find the product of (5+3i)(5-3i).... would this be 10? I think the 3i's cancel out..but i'm not sure?

OpenStudy (anonymous):

waiit.i just realized this is one where I foil..was thinking it was something else..

OpenStudy (anonymous):

so I actually might need help after all. I have (25)+(-15i)+(15i)+(-9i (squared)) Now what do I do with the squared -9i when I rewrite this problem?

OpenStudy (radar):

I think you may be dealing with the imaginary operator i. i is the sq root of -1\[\sqrt{-1}=i\]

OpenStudy (radar):

So in your answer 25-(-9)=34

OpenStudy (radar):

The product of +3i times -3i is :\[-9i ^{2}=(-9\times-1)=9\]